OpenStudy (anonymous):

Meg rowed her boat upstream a distance of 65 mi and then rowed back to the starting point. The total time of the trip was 18 hours. If the rate of the current was 4 mph, find the average speed of the boat relative to the water.

OpenStudy (anonymous):

let speed of boat be x time distance /velocity v up= x-4 v down= x+4 65/x-4 + 65/x+4 = 18 solve for x